OpenStudy (kuoministers):

Simplify 2^3 * 4n^2 * 2^-n HELP PLS :D

OpenStudy (kuoministers):

\[2^{3} \times 4n ^{2} \times 2^{-n}\]

OpenStudy (kuoministers):

\[2^{3} \times 2^{2}n ^{2} \times 2^{-n}\]

OpenStudy (kuoministers):

@Dido525

OpenStudy (kuoministers):

@RolyPoly

OpenStudy (kuoministers):

pls help

OpenStudy (kuoministers):

help @exraven

OpenStudy (kuoministers):

pls help anyone? @AravindG

OpenStudy (aravindg):